OpenStudy (kenialove):

At the same time a 32-foot water tower casts a 14-foot shadow , a nearby light posts a casts a 7-foot shadow. How tall is the light post? A. 7 ft B. 12 ft C. 16 ft D. 18 ft

OpenStudy (anonymous):

We want to use a proportion: 32/14 = x/7 32 * 7 = 14x 224 = 14x /14 on both sides x = 16

OpenStudy (kenialove):

Thanks so much